OpenStudy (anonymous):

A triangular prism has a volume of 64 yd3. If each dimension of the prism is reduced by half, what is the new volume of the prism? 32 yd3 16 yd3 128 yd3 8 yd3

OpenStudy (allank):

You need to use scale factors. Since all the dimensions are reduced by half, the linear scale factor (LSF) is 1/2. We want the change in volume, so calculate the volume scale factor (VSF) which is VSF = LSF^3 = 1/8 Thus new volume = 64*VSF

OpenStudy (anonymous):

so the answer is d

OpenStudy (allank):

Yes

OpenStudy (anonymous):

thnks

OpenStudy (allank):

no prob